OpenStudy (anonymous):

solve the proportion: x/2=5/15 x=

OpenStudy (anonymous):

awh put 2 on other side x= 5*2/15 =10/15

OpenStudy (anonymous):

x/2 = 5/15 (cross multiply) (15)(x) = (2)(5) 15x = 10 (divide both sides by 15 to get x by itself) 15x/15 = 10/15 x = 10/15 = 2/3